OpenStudy (anonymous):

Given A, B and C are sets, prove or disprove the following: (a) A subset B subset C <==> AxB subset BxC

OpenStudy (anonymous):

\[A \subseteq B \subseteq t C <=> A \times B \subseteq B \times C\]

OpenStudy (anonymous):

typing error..no t in this expression

OpenStudy (anonymous):

can probably do this working directly from the definitions

OpenStudy (anonymous):

first show \[A \subseteq B \subseteq C \implies A \times B \subseteq B \times C\]

OpenStudy (anonymous):

if and only if

OpenStudy (anonymous):

yeah you have to go both ways. first the implication above. you assume \[A \subseteq B \subseteq C \] and show \[A \times B \subseteq B \times C\]

OpenStudy (anonymous):

by definition \[A\times B=\{(a,b)|:a\in A, b\in B\}\] let \((a,b)\in A\times B\) then since \(A \subseteq B, a\in B\) and since \(B \subseteq C, b\in C \) therefore \((a,b)\in B\times C\)

OpenStudy (anonymous):

what about left side to right side ? how to prove it..